# Principled priors for Bayesian inference of circular models

### v2

@@ 7 Conclusion and Outlook
A principled framework for selecting priors for parameters of circular distributions is essential for constructing robust and accurate Bayesian circular models to prevent overfitting, especially when the model complexity increases. The framework proposed in this paper addresses the need to evaluate a prior’s ability to prevent overfitting, providing weakly informative priors that support the systematic construction of robust Bayesian circular models. Additionally, it offers a straightforward approach for selecting prior hyperparameters.
The derived priors are inherently designed to prevent overfitting and remain robust choices even when sufficient information is available. However, they may not always be the most appropriate choice in such cases, as more informative priors tailored to specific scenarios could perform better. In these situations, the proposed framework can be employed to rigorously evaluate model complexity and ensure the appropriateness of alternative priors.
With prior selection strategies addressed in this work, the next step involves developing general and intuitive procedures for constructing a variety of Bayesian circular models, including those with circular objectives, circular covariates, and circular responses. Future efforts will also focus on extending the framework to accommodate joint priors and providing conditional PC priors, which are vital for handling multi-parameter and multi-modal distributions and models.
@@ A.2.1 With Circular Uniform Base Model
[…]
@@
[…]
@@ A.3 The PC Prior for the Concentration Parameter of Wrapped Cauchy Distribution
[…]

### v3

@@ 6 Conclusion and Outlook
A principled framework for selecting priors for parameters of circular distributions is essential for constructing robust Bayesian circular models and for avoiding overfitting as model complexity increases. The framework proposed in this paper focuses on assessing whether a prior can properly guard against overfitting, thereby supporting a more reliable construction of Bayesian circular models. Additionally, it provides a simple and practical way to choose prior hyperparameters so that the resulting priors can be either strongly or weakly informative, depending on the level of prior knowledge available.
The derived priors are inherently designed to prevent overfitting and remain robust choices even when sufficient information is available. However, they may not always be the most appropriate choice in such cases, as more informative priors tailored to specific scenarios could perform better. In these situations, the proposed framework can be employed to rigorously evaluate model complexity and ensure the appropriateness of alternative priors. It is worth noting that the present analysis does not include a direct numerical diagnostic for prior overfitting. Instead, overfitting is assessed theoretically by examining whether a prior places sufficient mass near the chosen base model, which is clearly observed when representing the prior on the distance scale. While a dedicated quantitative metric could further support this assessment, introducing such tools is beyond the scope of this work and may be explored in future research.
With prior selection strategies addressed in this work, the next step involves developing general and intuitive procedures for constructing a variety of Bayesian circular models, including those with circular covariates, circular responses, or both. Future efforts will also focus on extending the framework to accommodate joint priors and providing conditional PC priors, which are vital for handling multi-parameter and multi-modal distributions and models.
Supplementary materials
The supplementary materials include the appendix for the derivation of the PC priors, and the code for simulation studies (in ibex cluster) and application (in R). They are available at https://github.com/XiangYEstats/PCpriors-circular.
@@ A.2.1 With Circular Uniform Base Model
$\int_{0}^{\frac{1}{2}}\pi\left(\ell\right)d\ell=1-\exp\left\{-\lambda\sqrt{1-\log\left(2\right)}\right\}.$ (47)
@@
$F\left(\ell\right)$ $=\frac{\int\pi\left(\ell\right)d\ell}{\int_{0}^{\frac{1}{2}}\pi\left(\ell\right)d\ell}=\frac{1-\exp\left\{-\lambda\sqrt{1-\sqrt{1-4\ell^{2}}+\log\left(\ell\right)+\frac{1}{2}\log\left(\frac{1+\sqrt{1-4\ell^{2}}}{1-\sqrt{1-4\ell^{2}}}\right)}\right\}}{1-\exp\left\{-\lambda\sqrt{1-\log\left(2\right)}\right\}}.$ (48)
@@ A.3 The PC Prior for the Concentration Parameter of Wrapped Cauchy Distribution
Since the normalizing constant is $\int_{0}^{1}\pi\left(\rho\right)d\rho=1$, the density remain the same.